OpenStudy (anonymous):

suppose a 1000$ was invested in an account with interest compounded continously. If at the end of 5 years the account holds 2000$. what was the interest rate on the account ? i just dont know how to solve for r

OpenStudy (kropot72):

I have already shown you the method. See your previous question.

OpenStudy (anonymous):

but how do i replace for r ?

OpenStudy (anonymous):

this is compounded continuously...

OpenStudy (kropot72):

The formula for continuous compounding is \[A=P \times e ^{rt}\] \[2000=1000\times e ^{5r}\] Taking logs of both sides: \[\ln 2000=\ln 1000+5r\] \[5r=\ln 2000-\ln 1000\] \[r=\frac{\ln 2000-\ln 1000}{5}\]

OpenStudy (anonymous):

0.06 ?

OpenStudy (anonymous):

or 0.14?

OpenStudy (kropot72):

\[r=\frac{\ln 2000-\ln 1000}{5}=0.1386\]

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)